# MoM: Linear Sequence Modeling with Mixture-of-Memories

### v1

@@ Abstract
† interns at Shanghai AI Laboratory; 🖂  corresponding authors (sunweigao@outlook.com, chengyu@cse.cuhk.edu.hk).
@@ 2 Preliminary
Softmax Attention
Given input $X=[x_{1},...,x_{n}]^{T}\in\mathbb{R}^{n\times d}$,where n is the sequence length and $x_{t}\in\mathbb{R}^{d}$ is the $t$-th input vector with $d$ dimensions. Transformer Vaswani (2017) softmax attention computes:
$\bm{Q}=\bm{X}\bm{W}_{Q},\bm{K}=\bm{X}\bm{W}_{K},\bm{V}=\bm{X}\bm{W}_{V},$ (1)
$\bm{O}=\mathrm{softmax}(\frac{\bm{Q}\bm{K}^{T}}{\sqrt{d}}\odot\bm{A})\bm{V},$ (2)
where $\bm{Q},\bm{K},\bm{V}\in\mathbb{R}^{n\times d}$, $\bm{A}\in\mathbb{R}^{d\times d}$ is an attention mask, $\odot$ denotes element-wise production.
The training time complexity is $O(n^{2})$ due to the $\bm{Q}\bm{K}^{T}$ operation. During inference, Transformers store all the $\bm{k},\bm{v}$ calculated as "$\bm{KV}$ cache" which can be viewed as memory states. As the $\bm{KV}$ cache grows with the input, the Transformers can be considered to separate the memory storage for each input token individually and have an unlimited memory capacity. This also results in a time complexity of $O(n)$ per token during inference.
@@
Figure 1: Framework of MoM. Each input token selectively activates and updates $K$ memory states, leaving non-activated memory states unchanged to avoid interference from current input. Additionally, we introduce a continuously activated shared memory. This figure presents the basic memory update mechanism; other mechanisms involving gating or more complex updates follow a similar approach.
@@ 3.2 MoM: Mixture-of-Memories
In our framework, the memory states function similarly to the experts in MoE. However, instead of relying on completely separate networks, these modules are individual RNN states embedded within a linear recurrent mechanism. This design allows for the isolation of memory updates while concurrently managing distinct types of information. It is important to note that MoM fundamentally differs from traditional MoE, as we will discuss in Appendix A. Figure 1 provides an overview of the MoM architecture. Below, we introduce the structure of the MoM layer and explain how this biologically-inspired architecture is implemented in the context of linear sequence modeling.
3.2.1 Router Network
@@ 3.2.2 Linear Recurrent Memory Module
After the router network, the input $\bm{x}_{t}$ is directed to top-$k$ linear recurrent modules, meaning that the top-$k$ memories are activated while the others remain inactive. For each activated memory module, indexed by $m$, we perform the following operation:
@@
GLA $\bm{M}_{t}=(\bm{a}_{t}^{T}\bm{1})\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$
Mamba2 $\bm{M}_{t}=\alpha_{t}\bm{M}_{t-1}+\beta_{t}\bm{k}_{t}^{T}\bm{v}_{t}$
DeltaNet $\bm{M}_{t}=(\bm{I}-\bm{k}_{t}^{T}\bm{k}_{t})\bm{M}_{t-1}+\beta_{t}\bm{k}_{t}^{T}\bm{v}_{t}$
G-DeltaNet $\bm{M}_{t}=\alpha_{t}(\bm{I}-\bm{k}_{t}^{T}\bm{k}_{t})\bm{M}_{t-1}+\beta_{t}\bm{k}_{t}^{T}\bm{v}_{t}$
TTT $\bm{M}_{t}=\bm{M}_{t-1}+\beta_{t}\nabla l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$
Titan $\bm{M}_{t}=\alpha_{t}\bm{M}_{t-1}+\beta_{t}\nabla l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$
Table 1: Memory Update Rules. We demonstrate that several current linear sequence models can be viewed as recurrent models in terms of memory updates, where $\alpha_{t},\beta_{t}\in(0,1)$ are data-dependent scaler, $\bm{a}_{t}$ is data-dependent vector, and $\gamma$ is a data-independent constant.
@@
340M Params 15B Tokens L=24, d=1024 Transformer++ 46.14 25.87 33.22 18.94 45.97 20.03 31.70
RetNet 5.90 9.28 22.41 6.91 40.05 18.59 17.19
HGRN2 11.53 17.34 24.08 12.67 43.84 17.35 21.14
GLA 11.26 16.78 27.85 12.77 43.90 17.68 21.71
GSA 6.36 16.87 21.90 14.60 42.18 16.72 19.77
Gated DeltaNet 20.53 23.24 28.55 14.98 44.91 16.48 24.78
MoM 30.79 26.05 29.63 13.84 44.79 20.41 27.59
1.3B Params 100B Tokens L=24, d=2048 Transformer++ 44.32 32.43 42.59 24.49 58.47 21.56 37.31
RetNet 13.62 22.59 33.46 15.43 53.79 19.79 26.45
HGRN2 12.35 23.24 33.19 19.10 55.27 19.65 27.13
GLA 27.61 30.93 35.04 22.27 56.28 19.45 31.93
GSA 23.25 32.80 35.57 22.96 57.05 20.65 32.05
Gated DeltaNet 30.25 27.65 34.06 23.22 58.23 20.36 32.30
MoM 41.14 34.30 37.08 24.11 58.59 21.03 36.04
@@ 4.1 Experimental Setups
Training.   We follow the training procedure described by Yang et al. (2023), utilizing the SlimPajama dataset Soboleva et al. (2023) sampled with 100B tokens and tokenized using the Mistral tokenizer Jiang et al. (2023). We train models from scratch with parameter sizes of 340M and 1.3B, respectively. For the 340M models, we train on 15B tokens with a batch size of 0.5M tokens. The warmup tokens count is set to 0.25M. We set the hidden ratio of our model to 3 to keep the overall parameter count approximately the same. For the 1.3B models, we train on 100B tokens with a batch size of 2M tokens. We utilized publicly available pretrained weights from Zhang et al. (2024) with exactly same configuration. The warmup tokens count is 1B. We employ AdamW optimizer Loshchilov et al. (2017); Sun et al. (2024b) with learning rate of 3e-4 with cosine learning rate schedule Zhou et al. (2020). The weight decay is set to 0.01 and gradient clipping is 1.0.
@@ 4.2.3 Long Context Tasks
Table 4: LongBench Results. All evaluations were done using Contributors (2023). Note: Sum = Summarization, FS = Few-shot, Syn = Synthetic.
@@
GLA expanded 15.08 20.15 28.28 13.30 41.65 18.74 22.87
GLA MoM 9.90 21.65 29.36 14.16 45.20 20.89 23.53
Gated DeltaNet expanded 18.26 24.27 30.03 17.74 48.34 19.26 26.32
Gated DeltaNet MoM 28.07 29.15 29.73 15.24 45.50 20.41 28.02
Table 5: Comparison Between Mixed Memory and Single Memory. We constructed MoM models using different memory update mechanisms. Separate memory segments yielded better performance compared to simply increasing the memory capacity of a single memory.
@@
To validate the effectiveness of our mixed memory mechanism, we compare our MoM model with mixed memories to a baseline model that uses an expanded single memory with the same activated memory capacity. We adopt the same memory update method as existing linear models and extend it within our MoM framework. For comparison, we adjust the dimensions of projected $\bm{k}$ and $\bm{v}$ in the original model to match the total size of all activated memories in the MoM model. Note that while our MoM model activates multiple memories during retrieval, the mixing process effectively consolidates these into a memory capacity equivalent to a single memory. In contrast, the expanded single memory approach increases the entire retrieval capacity to encompass the size of multiple activated memories. We evaluate its performance on the recall-intensive tasks in Table 5.
@@ 4.2.5 Efficiency
Figure 2: Efficiency of MoM. We demonstrate the inference time and GPU memory consumption required to generate 1K tokens at specific sequence lengths.
4.2.6 Training Loss Comparison
4.2.7 Ablation
@@
5 Related Work
@@
6 Conclusion
Appendix A Comparison between MoM and MoE
@@
Appendix B Dataset and Benchmark
@@
Appendix C Experiments Details
@@
Table 7: Complete Results of LongBench. (SQA: Single-doc QA, MQA: Multi-doc QA, Sum: Summarization, FS: Few-shot learning, Syn: Synthetic)
@@
Table 8: Complete Ablation Study Results. The complete results of the ablation studies on recall-intensive tasks.
@@
Table 9: Complete Ablation Study Results. The complete results of the ablation studies on recall-intensive tasks.

### v2

@@ Abstract
† interns at Shanghai AI Laboratory; 🖂  corresponding authors (sunweigao@outlook.com, chengyu@cse.cuhk.edu.hk). Weigao Sun is the project lead.
@@ 2 Preliminary
Figure 1: MoM Architecture. Each input token selectively activates and updates $K$ memory states, leaving non-activated memory states unchanged to avoid interference from current input. Additionally, we introduce a continuously activated shared memory. This figure presents the basic memory update mechanism; other mechanisms involving gating or more complex updates follow a similar approach.
@@ 3.2 MoM: Mixture-of-Memories
In our framework, the memory states function similarly to the experts in MoE. However, instead of relying on completely separate networks, these modules are individual RNN states embedded within a linear recurrent mechanism. This design allows for the isolation of memory updates while concurrently managing distinct types of information. It is important to note that MoM fundamentally differs from traditional MoE, as we will discuss in Appendix B. Figure 1 provides an overview of the MoM architecture. Below, we introduce the structure of the MoM layer and explain how this biologically-inspired architecture is implemented in the context of linear sequence modeling.
3.2.1 Router
@@ 3.2.2 Linear Recurrent Memory Module
After the router network, the input $\bm{x}_{t}$ is directed to top-$k$ linear recurrent modules, meaning that the top-$k$ memories are activated while the others remain inactive.
Each Memory. For each activated memory, indexed by $m$, we perform the following operation:
@@
GLA $\bm{M}_{t}=(\bm{a}_{t}^{T}\bm{1})\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$
DeltaNet $\bm{M}_{t}=(\bm{I}-\bm{k}_{t}^{T}\bm{k}_{t})\bm{M}_{t-1}+b_{t}\bm{k}_{t}^{T}\bm{v}_{t}$
G-DeltaNet $\bm{M}_{t}=a_{t}(\bm{I}-\bm{k}_{t}^{T}\bm{k}_{t})\bm{M}_{t-1}+b_{t}\bm{k}_{t}^{T}\bm{v}_{t}$
TTT $\bm{M}_{t}=\bm{M}_{t-1}+b_{t}\nabla l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$
Titans* $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+b_{t}\nabla l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$
Mamba2 $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+b_{t}\bm{k}_{t}^{T}\bm{v}_{t}$
@@
RWKV6 $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+\bm{k}_{t}^{T}\bm{v}_{t}$
RWKV7 $\bm{M}_{t}=a_{t}\bm{M}_{t-1}+b_{t}\nabla l(\bm{M}_{t-1};\bm{k}_{t},\bm{v}_{t})$
    For Titans, we have omitted details like momentum and norm optimization during gradient descent to keep the expression simplified.
Table 1: Memory Update Rules. We demonstrate that several current linear sequence models can be viewed as recurrent models in terms of memory updates, where $a_{t},b_{t}\in(0,1)$ are data-dependent scaler, $\bm{a}_{t}$ is data-dependent vector, and $\gamma$ is a data-independent constant.
@@
Avg.
(no FDA)
340M Params 15B Tokens L=24, d=1024 Transformer++ 46.14 25.87 33.22 18.94 45.97 20.03 31.70 28.81
RetNet 5.90 9.28 22.41 6.91 40.05 18.59 17.19 19.45
HGRN2 11.53 17.34 24.08 12.67 43.84 17.35 21.14 23.06
GLA 11.26 16.78 27.85 12.77 43.90 17.68 21.71 23.80
GSA 6.36 16.87 21.90 14.60 42.18 16.72 19.77 22.45
Gated DeltaNet 20.53 23.24 28.55 14.98 44.91 16.48 24.78 25.63
MoM 22.98 29.90 29.69 16.60 48.82 20.99 28.16 29.20
1.3B Params 100B Tokens L=24, d=2048 Transformer++ 44.32 32.43 42.59 24.49 58.47 21.56 37.31 35.91
RetNet 13.62 22.59 33.46 15.43 53.79 19.79 26.45 29.01
HGRN2 12.35 23.24 33.19 19.10 55.27 19.65 27.13 30.09
GLA 27.61 30.93 35.04 22.27 56.28 19.45 31.93 32.79
GSA 23.25 32.80 35.57 22.96 57.05 20.65 32.05 33.81
Gated DeltaNet 30.25 27.65 34.06 23.22 58.23 20.36 32.30 32.70
MoM 41.14 34.30 37.08 24.11 58.59 21.03 36.04 35.02
@@ 4.1 Experimental Setups
Training.   We follow the training procedure described by Yang et al. (2023), utilizing the SlimPajama dataset Soboleva et al. (2023) sampled with 100B tokens and tokenized using the Mistral tokenizer Jiang et al. (2023). We train models from scratch with parameter sizes of 340M and 1.3B, respectively. For the 340M models, we train on 15B tokens with a batch size of 0.5M tokens. The warmup tokens count is set to 0.25M. We set the hidden ratio of our model to 3 to keep the activated parameter count approximately the same. For the 1.3B models, we train on 100B tokens with a batch size of 2M tokens. We utilized publicly available pretrained weights from Zhang et al. (2024) with exactly same configuration. The warmup tokens count is 1B. We employ AdamW optimizer Loshchilov et al. (2017); Sun et al. (2024b) with learning rate of 3e-4 with cosine learning rate schedule Zhou et al. (2020). The weight decay is set to 0.01 and gradient clipping is 1.0.
Parameter Explanation. The parameter counts differ across models due to the varying modules they contain. For reference, the "340M" model refers to a configuration with 24 layers, each having a hidden size of 1024. Similarly, "1.3B" refers to 24 layers with a hidden size of 2048. Actual parameter variations may occur due to differences in specific linear modules. The primary goal of MoM is to enable sparse activation and expand the memory capacity of linear sequence models. To achieve this, we applied sparse activation only to the key and value projections, resulting in a minor increase in activated parameters. This small increase is well justified given the substantial improvements in performance. We’ve also conducted comparisons with models that traditionally use larger parameter counts to expand memory capacity. A detailed discussion on fairness is provided in Appendix F.
@@ 4.2.2 Commonsense Reasoning Tasks
Table 4: LongBench Benchmark Results. All evaluations were done using OpenCompass. Note: Sum = Summarization, FS = Few-shot, Syn = Synthetic.
@@
Transformer++ 46.14 25.87 33.22 18.94 45.97 20.03 31.70
MoM 22.98 29.90 29.69 16.60 48.82 20.99 28.16
MoM Hybrid 58.13 44.05 35.71 20.18 48.10 20.60 37.80
Model
ARC-e
$\uparrow$
ARC-c
$\mathrm{acc_{n}}$$\uparrow$
Hella.
$\mathrm{acc_{n}}$$\uparrow$
Lamb.
$\mathrm{acc}$$\uparrow$
PIQA
$\mathrm{acc}$$\uparrow$
Wino.
$\mathrm{acc}$$\uparrow$
Avg.
Transformer++ 44.91 25.94 34.95 26.90 64.31 51.07 41.35
MoM 44.65 24.74 36.54 27.93 66.16 51.78 41.97
MoM Hybrid 46.55 24.49 36.45 28.86 65.51 52.41 42.38
Table 5: Hybrid Model Performance. The hybrid model integrates 1 Transformer layer after every 7 MoM layers, resulting in only 3 Transformer layers across a total of 24 layers.
4.2.3 Long Context Tasks
Model Params
ARC-e
$\uparrow$
ARC-c
$\mathrm{acc_{n}}$$\uparrow$
Hella.
$\mathrm{acc_{n}}$$\uparrow$
Lamb.
$\mathrm{acc}$$\uparrow$
PIQA
$\mathrm{acc}$$\uparrow$
Wino.
$\mathrm{acc}$$\uparrow$
Avg.
GLA expanded 425M 42.34 22.95 34.56 20.45 63.00 50.12 38.90
GLA MoM 395M 42.85 24.15 36.60 23.23 63.22 49.88 39.99
Gated DeltaNet expanded 550M 43.60 24.66 37.80 26.90 64.47 50.51 41.32
Gated DeltaNet MoM  [^4] M 44.65 24.74 36.54 27.93 66.16 51.78 41.97
Model Params FDA SWDE SQUAD NQ TriviaQA Drop Avg.
GLA expanded 425M 15.08 20.15 28.28 13.30 41.65 18.74 22.87
GLA MoM 395M 9.90 21.65 29.36 14.16 45.20 20.89 23.53
Gated DeltaNet expanded 550M 18.26 24.27 30.03 17.74 48.34 19.26 26.32
Gated DeltaNet MoM  [^4] M 22.98 29.90 29.69 16.60 48.82 20.99 28.16
Table 6: Comparison of Mixture of Memories and Single Memory Expanded. We constructed MoM models using different memory update mechanisms. Separate memory segments yielded better performance compared to simply increasing the memory capacity of a single memory.
4.2.4 The Hybrid of MoM and Transformer
We delve deeper into the hybridization of MoM and Transformer layers by integrating 1 Transformer layer after every 7 MoM layers, resulting in only 3 Transformer layers across a total of 24 layers. The performance on commonsense reasoning and recall-intensive tasks is presented in the table 5. MoM-Hybrid demonstrates significantly improved results compared to Transformer models, despite using only 3 layers of global attention.
@@ 4.2.4 The Hybrid of MoM and Transformer
To validate the effectiveness of our mixed memory mechanism, we compare our MoM model with mixed memories to a baseline model that uses an expanded single memory with the same activated memory capacity. We adopt the same memory update method as existing linear models and extend it within our MoM framework. For comparison, we employed the commonly used method of expanding the single memory by expanding the dimension of $v$ to match the total size of all activated memories in the MoM model. Note that while our MoM model activates multiple memories during retrieval, the mixing process effectively consolidates these into a memory capacity equivalent to a single memory. In contrast, the expanded single memory approach increases the entire retrieval capacity which leads to a significant increase in the number of parameters. We evaluate their performance on common-sense reasoning tasks and recall-intensive tasks in Table 6.
@@
Figure 2: Inference Efficiency of MoM. We demonstrate the inference (decoding) time and GPU memory consumption required to generate 1K tokens at specific sequence lengths.
@@ 4.2.6 Efficiency
4.2.7 Length Extrapolation
We pretrained the models on the Slimpajama dataset with a 2K context length and conducted extrapolation experiments on various lengths using the FinewebPenedo et al. (2024) dataset. We extended the length to 32K to calculate perplexity (ppl). As shown in Fig 3, the Transformer model experienced a significant increase in ppl due to its poor extrapolation capability. Among the linear models, MoM achieved the best results.
Figure 3: Length Extrapolation. We extrapolated models trained on 2K sequences to a length of 32K for perplexity (ppl) evaluation. The ppl of Transformer++ rises sharply from 13.4 at 2k to beyond the display range, reaching 2271.3 at 32k.
Figure 4: Memory Load Balance Analysis. Token Routing Distribution Across Layers and Memories with and without Aux Loss.
4.2.8 Memory Analysis
Memory Load Balance Analysis. To evaluate whether each memory segment in MoM is effectively balanced during inference on downstream tasks, we analyzed the number of tokens routed to each layer using around 300k tokens from the ARC-easy benchmark. We visualized the results with and without auxiliary loss in Fig 4 with heatmaps. Due to the adoption of auxiliary loss, the memory segments in each layer are almost uniformly routed and activated.
Functions of Different Memories. We analyzed the memory routing within the model and identified the categories of tokens processed by each of the four memories. By examining an intermediate layer, we observed that, unlike traditional MoE in FFN layers where experts often lack clear specialization and focus on syntactic punctuationJiang et al. (2023), MoM displays distinct specialization among its memories. This finding suggests potential for exploring an MoE architecture where each memory serves a specific role. These distinctions among different memories are presented in Table 7.
4.2.9 Ablation Study
@@
Memory Index Type of Tokens Potential Function
1 Basic nouns/verbs/prepositions Simplify semantic information/causal logic
2 Proper nouns/scientific terms Specialized knowledge
3 Technical terms/adjectives Detailed memory
4 Interrogative words/incomplete nouns Fragmented/open information
Table 7: Functions of Different Memories. By analyzing the types of tokens routed to the model’s intermediate layer, we observed a degree of specialization among the different memories.
@@ 4 Interrogative words/incomplete nouns Fragmented/open information
5 Conclusion
6 Acknowledgments
This work is supported by the Shanghai Artificial Intelligence Laboratory.
Appendix A Related Work
@@
Appendix B Comparison between MoM and MoE
@@
Appendix C Dataset and Benchmark
@@
Appendix D Experiments Details
@@
Appendix E Training Loss Comparison
Appendix F Fairness
To enhance memory capacity, MoM applies sparse activation to the key and value projections. Although these projections constitute a small portion of the overall model parameters, this inevitably increases the parameter count. Due to differing linear model structures, aligning both parameter count and memory capacity exactly is challenging. Thus, to ensure fairness, we conduct comparisons from two perspectives: equal activated parameter count and equal memory capacity.
F.1 Equal Activated Parameter Count
To ensure a fair comparison of parameters, we reduced the MLP hidden ratio to 2 and retrained the MoM model using the same training configurations as in Section 4.1. Both MoM and Gated Deltanet were set with 400M activated parameters. Although the smaller hidden ratio might impact the model’s commonsense knowledge, we tested on commonsense reasoning tasks and recall-intensive benchmarks. MoM consistently outperformed Gated Deltanet in both tests, further validating the effectiveness of the MoM approach. The results are presented in Table 9
Model Params
ARC-e
$\uparrow$
ARC-c
$\mathrm{acc_{n}}$$\uparrow$
Hella.
$\mathrm{acc_{n}}$$\uparrow$
Lamb.
$\mathrm{acc}$$\uparrow$
PIQA
$\mathrm{acc}$$\uparrow$
Wino.
$\mathrm{acc}$$\uparrow$
Avg.
Gated DeltaNet 400M 46.04 23.55 35.18 27.01 66.05 50.83 41.44
MoM 400M 47.10 23.72 35.43 26.88 64.64 51.22 41.50
Model Params FDA SWDE SQUAD NQ TriviaQA Drop Avg.
Gated DeltaNet 400M 20.53 23.24 28.55 14.98 44.91 16.48 24.78
MoM 400M 24.16 25.59 29.46 15.36 46.15 18.35 26.51
Table 9: Comparison with the Same Activated Parameters. MoM and Gated DeltaNet with 400M activated parameters are tested.
F.2 Equal Memory Capacity
To ensure a fair comparison of memory capacity, we also compared the single extended memory model with the MoM model. Notably, the single extended memory model has more parameters than the activated parameters in the MoM due to the extension of the $v$ dimension. MoM expands memory more elegantly and significantly outperforms in both recall and commonsense tasks. This comparison result is presented in Table 6.
@@ F.2 Equal Memory Capacity
Table 10: Complete Results of LongBench. SQA: Single-doc QA, MQA: Multi-doc QA, Sum: Summarization, FS: Few-shot learning, Syn: Synthetic
@@
Table 11: Complete Ablation Study Results. The complete results on recall-intensive tasks.
@@
Table 12: Complete Ablation Study Results. The complete results on common-sense reasoning tasks.